3. What is the weight of your 20 kilogram backpack in SI, CGS, and US Customary units? Please state all assumptions.
Answer #1

Assumption - weight is measured at sea level where g is constant at 9.81 m/s2

Mass of backpack = 20 kg

Weight = mass x acceleration of gravity

In SI units, g = 9.81 m/s2

W = 20 kg x 9.81 m/s2

= (196.2 kg-m/s2) x (1N / kg-m/s2)

= 196.2 N

In CGS

Mass = 20 kg x 1000g/kg = 20000 g

g = 9.81 m/s2 x 100cm/m = 981 cm/s2

W = 20000 g x 981 cm/s2

= (19620000 g-cm/s2) x (1dyne/ g-cm/s2)

= 19620000 dyne

In US customary units

Mass = 20 kg x 2.205lbm/kg = 44.1 lbm

g = 9.81 m/s2 x 3.2808ft/m = 32.185 ft/s2

W = 44.1 lbm x 32.185 ft/s2

= 1419.12 lbm-ft/s2 x (1 lbf / 32.174 lbm-ft/s2)

= 44.1 lbf